### v2

# How Do Social Bots Participate in Misinformation Spread? A Comprehensive Dataset and Analysis

Abstract
The social media platform is an ideal medium to spread misinformation, where social bots might accelerate the spread. This paper is the first to explore the interplay between social bots and misinformation on the Sina Weibo platform. We construct a large-scale dataset that contains annotations of misinformation and social bots. From the misinformation perspective, this dataset is multimodal, containing 11,393 pieces of misinformation and 16,416 pieces of real information. From the social bot perspective, this dataset contains 65,749 social bots and 345,886 genuine accounts, where we propose a weak-supervised annotator to annotate automatically. Extensive experiments prove that the dataset is the most comprehensive, misinformation and real information are distinguishable, and social bots have high annotation quality. Further analysis illustrates that: (i) social bots are deeply involved in information spread; (ii) misinformation with the same topics has similar content, providing the basis of echo chambers, and social bots amplify this phenomenon; and (iii) social bots generate similar content aiming to manipulate public opinions.
How Do Social Bots Participate in Misinformation Spread?
A Comprehensive Dataset and Analysis
Herun Wan1        Minnan Luo1        Zihan Ma1 Guang Dai2        Xiang Zhao3 1Xi’an Jiaotong University    2SGIT AI    3National University of Defense Technology wanherun@stu.xjtu.edu.cn
1 Introduction
Social media platforms, like $\mathbb{X}$ (Twitter) and Weibo, have become major information sources, and information spreads faster than traditional media. Due to the nature of such platforms, there have been attempts to disseminate misinformation, which could polarize society (Azzimonti and Fernandes, 2023) and impact the economy (Zhou et al., 2024). Meanwhile, besides attracting genuine users, the straightforward operation of social media also makes it an ideal breeding ground for fully automatic programs. Such programs, also known as social bots, could help publish various news and provide emergency help. While social bots can be considered a mere technological instrument, their malicious misuse could seriously compromise online safety. They are proven behind many online perils, including election interference (Ng et al., 2022) and hate speech propaganda (Stella et al., 2019). Social bots are natural message amplifiers (Caldarelli et al., 2020), increasing the risk of spreading misinformation (Huang et al., 2022). Namely, misinformation and social bots are two major factors endangering online security, and they would work together to amplify their impact as Figure 1 shows.
Figure 1: An example of social bots amplifying information spread. Social bots would publish similar content to amplify misinformation spread and manipulate public sentiment and stance, increasing harmfulness.
Researchers make efforts to fight the never-ending plague of misinformation and malicious social bots. They mainly propose automatic detectors to identify misinformation (Shu et al., 2019) and social bots (Yang et al., 2022). Meanwhile, researchers also explore how different types of content (Nan et al., 2021) or propagation patterns (Vosoughi et al., 2018) influence misinformation spread. From the social bot perspective, bot communities (Tan et al., 2023b) and bots’ repost behaviors (Elmas et al., 2022) have been investigated. While existing works have provided valuable insights into investigating misinformation and social bots, relatively little attention has been paid to the interplay between them.
In this paper, we aim to bridge the gap of existing works, investigating the interplay between misinformation and social bots. We construct a novel dataset named MisBot [^1] We have published MisBot in GitHub., which simultaneously contains information and annotations of misinformation and social bots (§2). Specifically, we first define the structure of MisBot. We then collect misinformation from Weibo’s official management center [^2] https://service.account.weibo.com/, being available for users who have logged in.. After that, we propose a simple but effective debiasing method to mitigate entity biases and collect real information from two credible sources. We finally employ a weak-supervised annotator to annotate users who participate in the information spread automatically. From the misinformation perspective, MisBot contains multiple modalities, including post content, comments, repost messages, images, and videos. MisBot includes 11,393 pieces of misinformation and 16,416 pieces of real information. From the social bot perspective, MisBot includes 952,955 users participating in the information spread, covering 65,749 annotated social bots and 345,886 genuine accounts. Extensive experiments (§3) prove that (i) MisBot is the most comprehensive and the only one with misinformation and social bot annotations, (ii) misinformation and real information are distinguishable, where a simply detector achieves 95.2% accuracy, and (iii) MisBot has high social bot annotation quality, where human evaluations prove it. Further analysis illustrates (§4) that (i) social bots are deeply involved in information spread, where 29.3% users who repost misinformation are social bots; (ii) misinformation with the same topics has similar content, providing the basis of echo chambers, and social bots amplify this phenomenon; and (iii) social bots generate similar content aiming to manipulate public opinions, including sentiments and stances.
2 MisBot Dataset
The collection process of MisBot consists of four components: (i) Data Structure, which defines the dataset structure; (ii) Misinformation Collection, which collects multiple modalities in misinformation; (iii) Real Information Collection, which collects real information without entity biases; and (iv) Weak-Supervised User Annotation, which trains a weak-supervised model to automatically annotate related accounts.
2.1 Data Structure
Unlike news platforms, users publish posts on Weibo, thus, we annotate user posts as misinformation or real information, where the main language is Chinese. From the online information perspective, each instance is represented as $A=\{s,\mathcal{G}_{\textit{repost}},\mathcal{G}_{\textit{comment}},I,V,y\}$. It contains textual content $s$, repost graph $\mathcal{G}_{\textit{repost}}$, comment graphs $\mathcal{G}_{\textit{comment}}$, images $I$, videos $V$, and corresponding label $y$. From the account perspective, each instance is represented as $U=\{F,T,y\}$. It contains the attribute set $F$, published posts $T$, and the corresponding label $y$. Meanwhile, if a user reposts, comments, or likes a post, we consider it to be participating in the spread of information. We provide example instances in MisBot in Appendix A.1.
2.2 Misinformation Collection
We collect posts flagged as misinformation from Weibo’s official management center [^3] We present the platform overview in Appendix A.2.. This platform presents posts containing misinformation judged by platform moderators or police. Besides, it provides a brief judgment to explain why the post is flagged as misinformation, which provides a basis for identifying topics of misinformation. We present an example in Appendix A.3. We have collected all the misinformation since this platform was established. Specifically, the misinformation collected was published between April 2018 and April 2024. We spent about 10 months collecting 11,393 pieces of misinformation.
2.3 Real Information Collection
Existing misinformation datasets generally suffer from potential data bias (Chen et al., 2023), especially entity biases (Zhu et al., 2022), which would influence models’ generalization ability to unseen data. Thus, we design an entity debiasing method to mitigate entity biases. We first employ a keyphrase extractor (Liang et al., 2021) to obtain key entities from each misinformation. After filtering [^4] We present the filter strategy in Appendix A.4 ., we obtain 1,961 entities. We finally query the key entities in trusted sources to get real information. To ensure the authenticity and diversity of real information, we collect real information from two sources:
  • Verified news media is an official news account certified by the Weibo platform, which contains a red “verified” symbol and a verified reason [^5] We provide verified news accounts in Appendix A.5..
  • Trends on the platform contains posts sparking a lot of discussion in a short period.
We discuss the credibility of these two sources and provide examples and detailed query methods in Appendices A.6 and A.7. We obtained 8,317 and 8,099 pieces of real information, respectively.
2.4 Weak-Supervised User Annotation
Manual annotation or crowd-sourcing is labor-intensive and not feasible with large-scale datasets. Meanwhile, to ensure the scalability of MisBot, we propose a weak supervision learning strategy, enabling the automatic annotation. The construction of the weak-supervised annotator contains three phases: (i) preprocessing phase; (ii) training phase; and (iii) inference phase.
Preprocessing Phase
This phase preprocesses the training dataset for the weak-supervised annotator. We first collected 100,000 random accounts. Due to the randomness, these accounts could represent the entire Weibo environment, ensuring the diversity of accounts. We employ crowd-sourcing to annotate them, where the human annotators are familiar with social media. We summarize the general criteria for identifying a social bot on Weibo and write a guideline document for human annotators based on the criteria. We provide the criteria and guideline document in Appendix A.8. To ensure the annotation quality, each human annotator is required to read the document before annotating. Each annotator should also annotate 20 standard accounts that are obviously social bots or genuine accounts. Annotators who achieve more than 80% accuracy on standard accounts are reliable. We ensure that each account is annotated by three reliable human annotators. We provide the number of human annotators, the salary of annotators, and the annotation cost in Appendix A.9. We employ major voting to obtain the final annotations in this phase.
We filter in active accounts [^6] We provide the filter strategy based on human annotators’ feedback in Appendix A.10. and focus on them in MisBot, where there are three main reasons:
  • We aim to explore the involvement of social bots in misinformation and real information spread, where inactive users hardly participate in information spread.
  • Annotators mainly rely on posts in users’ timelines to make judgments, whereas inactive accounts cannot provide enough information to obtain reliable annotations.
  • Mainstream social bot detectors analyze accounts’ posts to identify bots, and we follow this to construct an annotation model. We employ active users to ensure credibility.
We obtained 48,536 active accounts from the 100,000 accounts, of which 18,132 are social bots and 30,404 are genuine accounts.
Training Phase
Different machine bot detectors have their strengths and weaknesses in the face of multiple social bots (Sayyadiharikandeh et al., 2020). Thus, we propose to employ multiple detectors as experts and employ an ensemble strategy to obtain the final annotations. In this phase, we leverage the following detectors:
  • Feature-based detectors leverage feature engineering on user attributes and adopt classic machine learning algorithms to identify social bots. We employ various attributes: (i) numerical: follower count, following count, and status count; and (ii) categorical: verified, svip, account type, and svip level. We employ MLP layers, random forests, and Adaboost as detectors.
  • Content-based detectors leverage NLP technologies to analyze user-generated content, where we employ name, description, and posts. We employ encoder-based language models, including BERT (Devlin et al., 2019) and DeBERTa (He et al., 2021) to identify social bots.
  • Ensemble detectors combine user attributes and generated content to identify social bots.
We provide the descriptions and settings of experts in Appendix A.11. We create an 8:1:1 split for the users from the preprocessing phase as train, validation, and test sets to train each expert.
Dataset Modalities Statistics
Content Comment Repost Image Video User Post Image Video User Bots Human
Datasets for misinformation detection.
(Shu et al., 2020) 23,196 19,200 0 2,063,442 0 0
(Nan et al., 2021) 9,128 0 0 0 0 0
(Li et al., 2022) 700 0 700 0 0 0
(Qi et al., 2023) 3,654 0 3,654 3,654 0 0
(Hu et al., 2023) 14,700 14,700 0 0 0 0
(Li et al., 2024) 23,789 10,178 0 803,779 0 0
Datasets for social bot detection.
(Feng et al., 2021b) 0 0 0 229,580 6,589 5,237
(Feng et al., 2022) 0 0 0 1,000,000 139,943 860,057
(Shi et al., 2023) 0 0 0 410,199 2,748 7,451
Datasets for the interplay between misinformation and social bots.
MisBot 27,809 61,714 7,328 952,955 65,749 345,886
Table 1: Summary of our dataset and recent datasets for misinformation and social bots. We first check each dataset’s modality and then report the related statistics. The $\star$ denotes that the publisher does not provide the original data in the corresponding paper. Our dataset is the largest and the only one with misinformation and social bot annotations, containing 27,809 instances.
Inference Phase
This phase annotates an unseen account based on the predictions from multiple experts. To ensure annotation quality, we filter in the experts achieving 80% accuracy, which is the same standard as human annotators, on the validation set. After that, a conventional method to integrate multiple predictions is to employ majority voting or train an MLP classifier on the validation set (Bach et al., 2017; Feng et al., 2022). Since the likelihood from classifiers may not accurately reflect true probabilities (Guo et al., 2017), also known as miscalibrated, we calibrate the likelihoods before the ensemble. We employ temperature scaling and select the best temperature on the validation set [^7] We provide the temperature settings in Appendix A.12.. We finally average the calibrated likelihoods to obtain the final annotations. Among the 952,955 accounts that participate in information spread in MisBot, 411,635 are active, of which 65,749 are social bots and 345,886 are genuine accounts.
Figure 2: The joint distributions of three content consistency metrics for misinformation and real information. Misinformation and real information illustrate different distributions, especially in Text and Similarity.
3 Basic Analysis of MisBot
MisBot is the most comprehensive
We compare MisBot with the recent datasets for misinformation and social bots, illustrated in Table 1. MisBot is the only dataset simultaneously containing misinformation and social bot annotations. Meanwhile, from the misinformation perspective, MisBot contains the most complete multi-modal information, including textual content, user comments, repost messages, images, videos, and related users. MisBot is the largest and contains the richest visual modal data for misinformation.
Misinformation and real information in MisBot are distinguishable
We aim to explore the role of social bots in amplifying misinformation spread, which requires misinformation and real information to be distinguishable. Thus, we analyze whether misinformation and real information are distinguishable from two perspectives: data distribution and misinformation detector.
From the data distribution perspective, we first explore the differences in multi-modal content consistency between misinformation and real information. We employ three metrics [^8] We provide the calculation formula in Appendix B.1.: (i) Text, where the higher value means the text is more consistent with the global texts; (ii) Similarity, where the higher value means the text and image are more consistent; and (iii) Image, where the higher value means the image is more consistent with the global images. Figure 2 visualizes the joint distribution of these metrics. For Text, real information illustrates higher textual content consistency, where the majority lies at 0.6, while misinformation shows lower, where some even show negative values. We speculate that real information would follow a certain pattern, while misinformation would not. For Similarity, real information illustrates relatively more consistency than misinformation, where real information might publish text-related images and misinformation might not. Thus, we could conclude that misinformation and real information present different multi-modal content consistency, being distinguishable. To further capture the image differences between misinformation and real information, we present the distribution of image categories and sentiments in Figure 14 in Appendix B.2. It illustrates that misinformation and real information present distinct distributions [^9] We provide the experiment settings and detailed analysis in Appendix B.2..
Models Accuracy F1-score Precision Recall
Vanilla $95.2_{\pm 0.6}$ $92.3_{\pm 0.8}$ $93.7_{\pm 1.7}$ $91.0_{\pm 1.0}$
w/o Interaction $\mathop{81.6^{\star}_{\pm 4.5}}\limits_{14.2\%\downarrow}$ $\mathop{77.3^{\star}_{\pm 4.1}}\limits_{16.2\%\downarrow}$ $\mathop{64.4^{\star}_{\pm 5.9}}\limits_{31.3\%\downarrow}$ $\mathop{97.3^{\star}_{\pm 1.2}}\limits_{7.0\%\uparrow}$
w/o Vision $\mathop{94.1^{\star}_{\pm 0.5}}\limits_{1.1\%\downarrow}$ $\mathop{90.3^{\star}_{\pm 1.0}}\limits_{2.2\%\downarrow}$ $\mathop{94.2^{\dagger}_{\pm 1.2}}\limits_{0.4\%\uparrow}$ $\mathop{86.8^{\star}_{\pm 2.1}}\limits_{4.6\%\downarrow}$
w/o Extra $\mathop{78.5^{\star}_{\pm 5.2}}\limits_{17.5\%\downarrow}$ $\mathop{74.5^{\star}_{\pm 4.3}}\limits_{19.3\%\downarrow}$ $\mathop{60.6^{\star}_{\pm 6.0}}\limits_{35.4\%\downarrow}$ $\mathop{97.3^{\star}_{\pm 1.0}}\limits_{6.9\%\uparrow}$
Table 2: Performance of baseline and variants. We report the mean and standard deviation of ten-fold cross-validation. We also report the performance changes and conduct the paired t-test with vanilla, where $\star$ denotes the p-value is less than $0.0005$ and $\dagger$ denotes otherwise. Misinformation and real information are distinguishable with the help of user interactions.
From the misinformation detector perspective, we design a simple misinformation detector to verify whether the detector could identify misinformation in MisBot [^10] We provide the details of this model in Appendix B.3.. We present the performance of the detector and variants in Table 2. This simple detector achieves remarkable performance, where the f1-score reaches 92.3%. The ideal performance proves that misinformation and real information are easily distinguished by a machine detector, which helps explore the differences between social bots in spreading misinformation and real information. Meanwhile, the detector without interaction drops to 77.3% on f1-score, illustrating the effectiveness of user reactions, which coincides with our speculation that social bots might have different social patterns in misinformation and real information [^11] We provide more analysis about the ablation study in Appendix B.4..
MisBot has high social bot annotation quality, where the weak-supervised annotator is reliable
The construction of the weak-supervised annotator contains three phases, where the results from each phase are reliable:
  • Preprocessing phase. We recruited 315 human annotators, each of whom annotated 1,000 accounts and 20 standard accounts (the annotators did not know the standard accounts). Among them, 300 human annotators achieved more than 80% accuracy on the standard accounts. The average accuracy of the reliable annotators on the standard accounts is 93.75%. For the agreement between human annotators, the average Fleiss’ Kappa is 0.4281, showing moderate agreement.
  • Training phase. We employed multiple detectors aiming to identify various social bots. To ensure the weak-supervised annotator’s credibility, we filtered in detectors achieving 80% accuracy and obtained 4 detectors [^12] We provide the performance of each detector and the corresponding temperature in Appendix B.5.. The accuracy on the test set reaches 85.03%, which is higher than TwiBot-20 (Feng et al., 2021b) and TwiBot-22 (Feng et al., 2022), illustrating credibility.
  • Inference phase. To evaluate annotation quality in MisBot, we randomly sample 50 social bots and 50 genuine accounts and manually annotate them through a human expert. The Cohen’s Kappa between the human expert and automatic annotator is 0.74, showing good agreement.
4 Misinformation and Social Bots
Social bots are deeply involved in information spread
We check the bot percentage in MisBot:
  • The whole MisBot contains 952,955 accounts, of which 411,635 are active. There are 65,749 social bots, accounting for 15.97%.
  • Among 5,750 accounts publishing misinformation, there are 3,799 active accounts. There are 767 social bots, accounting for 20.19%.
  • Among 226,235 accounts participating in the misinformation spread, 95,360 are active. There are 13,020 social bots, accounting for 13.65%.
  • Among 749,763 accounts participating in the real information spread, 325,414 are active. There are 54,253 social bots, accounting for 16.67%.
Figure 3: Probability density distributions of the percentage of social bots in information reposting and commenting. Social bots are deeply involved in information reposting and commenting.
Figure 3 further presents the distribution of social bots in information reposting and commenting. The average bot percentage of misinformation reposting and commenting is 29.3% and 10.9%, respectively, while the percentage of real information reposting and commenting is 31.1% and 14.7%. It illustrates that the distribution of misinformation and real information is similar, with slightly more social bots participating in spreading real information than misinformation. Meanwhile, reposting tends to have a higher bot percentage than commenting. Thus, we could conclude that social bots are deeply involved in information spread, where the main spread method is to repost information.
Misinformation with the same topics has similar content, providing the basis of echo chambers, and social bots amplify this phenomenon
We first group all pieces of misinformation into clusters with the same topics according to the judgment [^13] We provide the clustering algorithm in Appendix C.1.. We group 11,393 pieces of misinformation into 2,270 clusters, each of which represents a specific topic or event, e.g., “The last two minutes of the air crash”. We aim to explore the textual content similarity of misinformation with the same topics and across different topics.
We first visualize the misinformation content representations in Figure 4, where we select 10 topics with the highest number of misinformation items. It shows the BERT representation of misinformation content using t-SNE dimensionality reduction [^14] We provide a detailed description of each selected topic in Appendix C.2.. The silhouette score is 0.29, illustrating that the clusters are cohesive. It indicates that each cluster shares similar content while different clusters share significant differences. It initially suggests that the misinformation environment is homogeneous, providing the basis for echo chambers.
Figure 4: Visualization of misinformation content representations within the top 10 topics. Each dot corresponds to a piece of misinformation, colored according to its topic. The topic labels annotated by the judgment are plotted at each cluster center. We also calculate the silhouette score (×$\times 100$). The cohesive clusters indicate misinformation about the same topic having similar content, providing the basis of echo chambers.
We conduct further numerical analysis by calculating semantic-level and token-level pairwise scores between any two clusters, where higher scores mean the content of the two clusters is more similar. For semantic level, we employ the cosine similarity of the BERT representations, while for token level, we leverage the ROUGE-L score [^15] We provide the experiment settings and detailed analysis in Appendix C.3.. Figure 16 in Appendix C.4 presents the heatmap of the pairwise scores. It illustrates that the values on the diagonal are much greater. For semantic level, the average diagonal value is 0.9448, and the others’ average is 0.5847. For token level, the average diagonal value is 0.781,5 and the others’ average is 0.0773. The numerical results emphasize that misinformation with the same topics has similar content, and misinformation with different topics has distinct content.
Then we explore how social bots participate in misinformation spread. Inspired by existing works, we consider an account a potential echo chamber member if it engages in at least two misinformation spreads in the same cluster. Figure 5 displays the distribution of bot percentage among echo chamber members and non-members within various clusters. It illustrates that around 18% non-members are social bots. Meanwhile, the members do not contain bots in about half of the clusters. However, in the other half, members exhibit a higher bot percentage across most clusters compared to non-members, reaching up to 50% in many clusters. We speculate that social bots might reinforce the echo chamber effect by actively engaging in misinformation spread with the same topics.
Figure 5: Bot percentage distribution comparison between echo chamber members and non-members across various clusters. Bot percentage among echo chamber members is generally higher than among non-members.
Social bots generate similar content aiming to manipulate public opinions
Online information consumers are reluctant to process information deliberately (Möller et al., 2020), becoming susceptible to cognitive biases (Pennycook et al., 2018; Vosoughi et al., 2018). We aim to explore how public opinion changes and how social bots potentially manipulate it. We focus on how public sentiments and stances change in MisBot [^16] We provide information on obtaining the sentiments and stances of a social text in Appendix C.5..
  • For public sentiments, we categorize sentiments into neutral and non-neutral (including happy, angry, surprised, sad, and fearful). Figure 17 in Appendix C.6 presents sentiment distribution in different social texts. It illustrates that misinformation would publish more emotional content, while real information would naturally be reported. On the other hand, public reactions are always emotional, where misinformation shows more anger while real information shows more happiness. Thus, public sentiments are emotional. We further explore the degree or extent to which public sentiments change over the information spread, introducing a variation measure [^17] We provide the detailed introduction of this measure in Appendix C.7.:
    $v_{\Delta}=\sum_{k=1}^{n}|f(x_{k})-f(x_{k-1})|,$
    where $f(x_{k})$ denotes neutral sentiment proportion at time $x_{k}$. Figure 6 visualizes sentiment variation distribution, where a larger value means a more drastic change. The average values of misinformation and real information are 0.287 and 0.225. It illustrates that public sentiment changes are dramatic during information spread.
    Figure 6: The distribution of neutral sentiment variation. Public sentiment changes are dramatic during information spread, with misinformation slightly more drastic.
    Figure 7: The proportion of comments with different stances as the comments increase. Public stances become increasingly polarized, where misinformation contains more comments with clear stances.
  • For public stances, we categorize stances into support, oppose, and neutral. Figure 7 presents the proportion of each stance with the comments increasing over time. A striking finding is that only about 1% accounts explicitly expressed a supportive stance, while the majority are neutral or opposed. Meanwhile, misinformation consistently presents higher opposition and lower neutrality. It illustrates that public stances become more polarized as the information spreads, where the neutral ratio suffers a drop of around 11%.
Therefore, we can conclude that as the information spreads, public opinions, including sentiment and stance, become polarized, especially regarding misinformation. We then numerically prove the correlation between polarization and social bots by the Pearson correlation coefficient [^18] We provide the calculation method and more analysis in Appendix C.8.: the number of social bots demonstrates strong correlations with the number of comments with non-neutral stances (r = 0.6661) and sentiments (r = 0.6750). The relatively high correlation coefficients indicate that social bots might influence public opinion.
Figure 8: Distribution of social bots’ and humans’ semantic similarities, where social bots present higher similarities. Namely, social bots would publish similar content to manipulate public opinions.
We further explore social bot characteristics in information spread. Figure 8 presents the semantic similarity of social bots and genuine accounts [^19] We provide the calculation method in Appendix C.9.. It illustrates that social bots present higher values. Namely, social bots would publicize similar content and amplify the bandwagon effect, where online users adopt behaviors or actions simply because others are doing so, influencing the information spread (Wang and Zhu, 2019). Figure 18 in Appendix C.10 illustrates the stances and sentiments of comments published by social bots. It demonstrates that social bots publish more emotional content and comments with clear stances. The results enhance the finding that social bots generate similar content, aiming to manipulate public opinions.
5 Related Work
5.1 Misinformation Detection
Mainstream detectors focuses on the information content, including text (Hartl and Kruschwitz, 2022; Xiao et al., 2024; Gong et al., 2024), images (Liu et al., 2023a; Zhang et al., 2024b, d), and videos (Tan et al., 2023a; Bu et al., 2024). They extract features such as emotion (Zhang et al., 2021) and employ neural networks such as graph neural networks (Tao et al., 2024; Zhang et al., 2024f; Lu et al., 2024) or neurosymbolic reasoning (Dong et al., 2024) to characterize information. Besides information content, the context, such as user interactions (Shu et al., 2019; Lu and Li, 2020), user profile (Sun et al., 2023; Xu et al., 2024), and evidence (Chen et al., 2024) provide helpful signals to detect misinformation. These models would model propagation patterns (Cui and Jia, 2024), construct news environments (Yin et al., 2024), or extract multi-hop facts (Zhang et al., 2024a) to enhance detection performance. Recently, to combat LLM-generated misinformation (Zhang et al., 2024e; Venkatraman et al., 2024), models employing LLMs (Wan et al., 2024; Nan et al., 2024) through prompting (Guan et al., 2024; Hu et al., 2024) and in-context learning (Wang et al., 2024) have been proposed.
5.2 Social Bot Detection
Social bot detectors fall into feature-, content-, and graph-based. Feature-based models conduct feature engineering for accounts (Feng et al., 2021a; Hays et al., 2023). Content-based models employ NLP techniques (Lei et al., 2023; Cai et al., 2024) to characterize the content. Graph-based models model user interactions as graph structures and employ graph neural networks (Feng et al., 2021c; Yang et al., 2023b; Zhou et al., 2023; Liu et al., 2024) in a semi-supervised way to identify bots. Many researchers are committed to exploring the risks and opportunities LLMs bring to bot detection (Tan and Jiang, 2023; Feng et al., 2024).
5.3 Social Media Safety
Social media safety has become more crucial (Mou et al., 2024), where misinformation and social bots are two main factors harming online safety. Numerous datasets for misinformation (Li et al., 2024; Qazi et al., 2024; Lin et al., 2024; Chen and Shu, 2024) and social bots (Feng et al., 2021b, 2022; Shi et al., 2023) are proposed. Based on these datasets, the generalization of detectors (Zhang et al., 2024c; Assenmacher et al., 2024), misinformation propagation pattern (Aghajari, 2023; Ashkinaze et al., 2024), how to mitigate misinformation spread (Konstantinou and Karapanos, 2023; Su et al., 2024; Ghosh et al., 2024), health-related misinformation (Yang et al., 2024; Shang et al., 2024), source credibility (Carragher et al., 2024; Mehta and Goldwasser, 2024), user profiling (Morales et al., 2023; Zeng et al., 2024), and bot communities (Liu et al., 2023b; Tan et al., 2023b) are investigated. However, relatively little attention has been paid to the interplay between misinformation and social bots, thus, we bridge the gap in this paper.
6 Conclusion
We explored the interplay between misinformation and social bots. We proposed a novel dataset named MisBot containing information and annotations of misinformation and social bots. MisBot is the most comprehensive, and misinformation and real information are distinguishable, and social bots have high annotation quality. The analysis illustrates that (i) social bots are deeply involved in information spread; (ii) misinformation provides the basis of echo chambers, and social bots amplify this phenomenon; (iii) social bots generate similar content aiming to manipulate public opinions.
Limitation
Our proposed dataset is the largest containing misinformation and social bot annotations simultaneously. Meanwhile, it contains multiple modalities of news, including images and video,s and interactions of users, including comments and reposts. However, due to the focus on news spread, it does not contain interactions like the friend relationship, missing potential relations between social bots and genuine accounts. Meanwhile, we propose a weak-supervised framework to annotate social bots, whose accuracy is similar to crowd-sourcing. However, it struggles to achieve better recall and might miss several social bots. Finally, the experiments in this work focus primarily on the Sina Weibo platform. We expect to expand our experiments and analysis to other social media platforms such as $\mathbb{X}$ (Twitter) or Reddit, in future work.
Ethics Statement
The research of misinformation and social bots is essential in countering online malicious content. This research demonstrates that social bots would amplify the spread of misinformation, enhancing echo chambers and manipulating public opinions. However, this work may increase the risk of dual-use, where malicious actors may develop advanced social bots to spread misinformation. We will establish controlled access to ensure that the data and trained model checkpoint are only publicly available to researchers. Meanwhile, we will hide the privacy information in the dataset when we publish it.
Our models are trained on crowd-sourced data, which might contain social biases, stereotypes, and spurious correlations. Thus, our model would provide incorrect annotations. We argue that the predictions of our models should be interpreted as an initial screening, while content moderation decisions should be made with experts in the loop.
Appendix A Details of MisBot Dataset
Figure 9: The examples in MisBot. We present (a) a misinformation example, (b) a real information example, and (c) a Weibo account example. We translate original information into English and conceal the private information.
A.1 Examples in MisBot
Formally, an online information instance is represented as $A=\{s,I,V,\mathcal{G}_{\textit{repost}},\mathcal{G}_{\textit{comment}},\mathcal{U},y\}$. The image set $I=\{I_{i}\}$ contains multiple images while the video set $V=\{V_{i}\}$ contains multiple videos. The repost graph $\mathcal{G}_{\textit{repost}}=\{\mathcal{V},\mathcal{E},\mathcal{T}\}$ is a dynamic text-attributed graph (or tree) where the center node is the information content and another node $v\in\mathcal{V}$ denotes a repost text, $e=(v_{i},v_{j})\in\mathcal{E}$ denotes a repost relation connecting $v_{i}$ and $v_{j}$, and $\mathcal{T}:\mathcal{V}\rightarrow\mathbb{R}$ denotes the published time of each node. $\mathcal{G}_{\textit{comment}}=\{\mathcal{G}_{\textit{comment}}^{i}\}$ denotes the comment graph set, where each comment graph $\mathcal{G}_{\textit{comment}}^{i}$ is a dynamic text-attributed graph (or tree). Each comment graph is similar to the repost graph except for the center node, where the center node is a comment that directly comments on the information. Besides, a Weibo account instance is represented as $U=\{F,T,y\}$. The feature set contains follower count, following count, status count, verified (2 types), svip (2 types), account type (10 types), and svip level (6 types). The post set $T$ contains the most recent five posts in the user timeline. We provide a piece of misinformation, real information, and a Weibo account example in Figure 9.
Figure 10: The overview of the Weibo’s official management center. We conceal private or unrelated information and translate the main information into English. We highlight the misinformation items.
A.2 Management Center
The Weibo’s official management center is a Weibo official. Here is the link: https://service.account.weibo.com/?type=5&status=0. If the users are logging into the platform for the first time, it will redirect to the Weibo homepage (https://weibo.com/). After logging in with a Weibo account, entering the platform again will lead to the right platform homepage. Figure 10 shows the overview of this platform, where we conceal private or unrelated information and translate the main language into English. If the users successfully log into this platform, they will view a similar website. It is worth noting that the number of instances that each logged-in user can access per day is limited, so it took us about 10 months to collect all the misinformation on the platform.
Post flagged as misinformation: Recently, in xxx, a "naughty child" took scissors and cut off the hair of a female customer in a barber shop when no one was paying attention. After the female customer called the police and negotiated, the parents compensated 11,500 yuan.
Judgment: After investigation, it was found that the Weibo post claiming that "a woman’s hair was cut off by a naughty child and her parents paid her 10,000 yuan in compensation" actually happened in May 2023, not recently. The respondent’s speech is "outdated information" and constitutes "publishing false information"
Table 3: An example of misinformation and corresponding judgment (translated into English). The judgment provides a basis for identifying misinformation topics.
A.3 Misinformation Example
After logging in to the platform, it mainly contains users’ posts flagged as misinformation and a corresponding judgment. The platform moderators or police flag the misinformation and publish the judgment. We provide an example in Table 3. The judgment is the same for different pieces of misinformation on the same event.
A.4 Entity Filter
We obtained 7,445 entities using the keyphrase extractor. We employ two strategies to filter out noisy entities:
  • Frequency less than 10. These entities appear occasionally in misinformation and are unlikely to cause entity bias. We only focus on common entities that appear in large numbers in misinformation, so we need to ensure that they appear at a similar frequency in real information.
  • The number of characters is 1. These entities might come from the noises of the keyphrase detector. Meanwhile, these entities may not contain enough semantic information.
After filtering, we obtained 1,961 entities. We believe these entities are common and contain rich semantic information. As a result, it would mitigate the effects of entity bias if real information also contains these entities.
A.5 Verified Accounts
We employ 8 verified news accounts, and Table 4 presents the information about them. They have a red “verified” symbol. When an account has more than 10,000 followers and this account has been read more than 10 million times in the last 30 days, it can obtain the red “verified” symbol.
Homepage Follower Count Status Count Discussion Count
https://weibo.com/u/1496814565 33.8 million 225.3 thousand 334.0 million
https://weibo.com/u/5044281310 32.6 million 163.2 thousand 573.0 million
https://weibo.com/u/1618051664 111.0 million 302.6 thousand 1.6 billion
https://weibo.com/u/1974808274 3.3 million 58.8 thousand 27.3 million
https://weibo.com/u/2028810631 107.0 million 166.4 thousand 469.0 million
https://weibo.com/u/2656274875 137.0 million 187.8 thousand 3.7 billion
https://weibo.com/u/1784473157 81.5 million 246.5 thousand 786.0 million
https://weibo.com/u/1642512402 62.4 million 224.4 thousand 410.0 million
Table 4: The information about the selected verified news accounts. We provide the homepage links of them. They have a huge number of followers and discussions.
Figure 11: The overview of the two search pages. The red box presents the search functions.
A.6 Source Credibility
Here we discuss the credibility of the two real information sources:
  • Verified news media. These accounts are operated by legitimate news media and verified by the Weibo platform. Thus we believe this source is credible.
  • Trends on the platform. Weibo is a responsible social platform, where content moderators are efficient. As a result, the content moderation mechanism makes it easier to moderate posts with a lot of discussion. Because users would report the posts that they think are fake. After receiving reports, moderators only need to verify the post content instead of the whole discussion. It takes only a few days to moderate misinformation on the training. Meanwhile, the posts we collected are from one month ago in the trend. There is plenty of time for moderation.
A.7 Query Method
We mainly employ the official search function of the Weibo platform to search the given entity. Given an entity, the search function will return several posts containing the entity.
  • Verified news media. After entering a specific account’s homepage, we could use the search function to search posts in this account.
  • Trends on the platform. Given an entity, such as happy, we collect posts in the trends using https://s.weibo.com/weibo?q=happy&xsort=hot.
Figure 11 presents an overview of these two search functions, where the red box illustrates the search function.
A.8 Annotation Guideline
We first summarize the general criteria to identify a social bot on Weibo: (i) reposting or publishing numerous advertisements, (ii) devoted fans of a star publishing numerous related content, (iii) containing numerous reposting content without pertinence and originality, (iv) publishing numerous unverified and negative information, (v) containing numerous posts with the “automatically” flags, (vi) repeated posts with the same content, and (vii) containing content that violates relevant laws and regulations.
Based on the criteria, we write a guideline document for human annotators in Figures 12 and 13. Each human annotator must read this document before annotating.
Figure 12: The overview of the guideline document, where we translate it into English. The human annotators are required to read this document before annotation.
Figure 13: The overview of the guideline document (cont.).
A.9 Annotation Cost
Each human annotator is required to annotate 1,000 accounts plus 20 standard accounts. If a human annotator achieves more than 80% accuracy on the standard accounts, we will adopt the annotator’s annotation. We will pay 200 yuan (about 28 dollars) for each qualified annotator. We recruited 315 annotators and, 300 are qualified. The crowd-sourcing takes about 60 days and costs 60,000 yuan.
A.10 Active Accounts
We focus on the active accounts in MisBot and this paper. According to the human annotators’ feedback and the characteristics of the Weibo platform. If an account publishes more than five posts with a length of no less than five characters in the timeline, then we consider this account active.
A.11 Expert Settings
In the training phase, we leverage three categories of social bot detectors as experts:
Feature-based Detectors
We first preprocess the selected initial features to obtain the features for classifiers. For the numerical features (including follower count, following count, and status count), we employ z-score normalization:
$z=\frac{x-\mu}{\sigma},$
where $x$ is the initial feature, $z$ is the preprocessed feature, and $\mu$ and $\sigma$ are the average and standard deviation in the training set. The average values are $5074.88$, $420.59$, and $1432.10$, while the standard deviation values are $283145.61$, $584.40$, and $1373.91$. For the categorical features (including verified, svip, account type, and svip level), we employ one-hot to obtain the initial representations. After that, we concatenate numerical and categorical representations to obtain the account representation $\boldsymbol{x}_{\textit{f}}$. After that, We employ MLP layers, random forests, and Adaboost as detectors. We adopt three feature-based experts (three classic classifiers).
Content-based Detectors
We employ name, description, and posts to identify social bots. We assuming the notation of name is $\boldsymbol{s}_{\textit{name}}$, of description is $\boldsymbol{s}_{\textit{desc}}$, and of posts is $\{\boldsymbol{s}_{\textit{post}}^{i}\}_{i=1}^{N}$ (here are $N$ posts). Given a text $\boldsymbol{s}$, we employ encoder-based language model to obtain the representation:
$\boldsymbol{x}=\mathrm{LM(\boldsymbol{s})}.$
For posts, we average the representation:
$\boldsymbol{x}_{\textit{post}}=\frac{1}{N}\boldsymbol{x}^{i}_{\textit{post}}.$
We feed the representations into an MLP layer to identify social bots. We employ the pre-trained parameters of the encoder-based language models and do not update the parameters. We employ the parameters in the Hugging Face for BERT [^20] Here is the model link. and DeBERTa [^21] Here is the model link.. We adopt six content-based experts (two encoder-based language models and three categories of texts).
Ensemble Detectors
We first employ MLP layers to transfer the feature-based and content-based representations and concatenate them:
$\boldsymbol{x}=\|_{i\in\{\textit{f},\textit{name},\textit{desc},\textit{post}\}}\mathrm{MLP}(\boldsymbol{x}_{i}).$
We adopt two ensemble experts (two encoder-based language models).
For all experts, we do not update the language model parameters. We set the hidden dim as 256, learning rate as $10^{-4}$, weight decay as $10^{-5}$, batch size as 64, dropout as 0.5, optimizer as Adam, activation function as LeakyReLU.
We do not employ graph-based detectors because neighbor information is hard to access on the Weibo platform and would cost a lot during the inference process. Besides, the automatic annotator already achieves acceptable annotation quality.
A.12 Temperature Settings
Temperature scaling is a post-precessing technique to make neural networks calibrated. It divides the logits (the output of the MLP layers and the input to the softmax function) by a learned scalar parameter,
$p_{i}=\frac{e^{z_{i}/\tau}}{\sum_{j\in\mathcal{Y}}e^{z_{j}/\tau}},$
where $\mathcal{Y}$ denotes the label set, $p_{i}$ is the probability of belonging to category $i$. We learn the temperature parameter $\tau$ on the validation set. We conduct a grid search from 0.5 to 1.5 with an interval of 0.001, obtaining the optimal value by minimizing the expected calibration error on the validation set.
Appendix B Details of Basic Analysis
B.1 Content Agreement Metrics
These three metrics are proposed to calculate the multi-modal content consistency, where a higher value means higher consistency. Formally, assuming each information instance is presented as $(T_{i},I_{i})$ (here we only focus on the textual and image content). Meanwhile, the information set is represented as $\{(T_{i},I_{i})\}_{i=1}^{N}$ (misinformation set or real information set). Given an instance $(T_{i},I_{i})$, we calculate Text as:
$\textit{text}_{i}=\frac{1}{N}\sum_{j=1}^{N}\mathrm{cosine}(\mathrm{BERT}(T_{i}),\mathrm{BERT}(T_{j})),$
where $\mathrm{cosine}(\cdot)$ denote the cosine similarity function, $\mathrm{BERT}(\cdot)$ denote the BERT encoder [^22] Here is the model link.. We calculate Similarity as:
$\textit{similarity}_{i}=\mathrm{cosine}(\mathrm{CLIP}(T_{i}),\mathrm{CLIP}(I_{i})),$
where $\mathrm{CLIP}(\cdot)$ denote the CLIP encoder [^23] Here is the model link.. We calculate Image as:
$\textit{image}_{i}=\frac{1}{N}\sum_{j=1}^{N}\mathrm{cosine}(\mathrm{ViT}(I_{i}),\mathrm{ViT}(I_{j})),$
where $\mathrm{ViT}(\cdot)$ denote the ViT encoder [^24] Here is the model link..
B.2 Image Distribution
To further explore the differences in image distribution between misinformation and real information, we check the categories of the image in information. We select four common categories: (i) person, (ii) emoji pack, (iii) landscape, and (iv) screenshot. We then investigate the sentiments of person and emoji pack, where person is realistic and emoji pack is virtual. The sentiments include neutral and non-neutral (angry, surprised, fearful, sad, and happy). To obtain the categories and sentiments, we employ pre-trained CLIP (Radford et al., 2021) [^25] Here is the model link. in zero-shot format. Figure 14 presents the image distribution of misinformation and real information. Images in real information tend to focus on people, while misinformation prefers to publish screenshots. Regarding sentiment, most of the images related to people in both real and misinformation are non-neutral, proving that information publishers tend to employ appealing pictures. For virtual images, emoji packs in real information are predominantly neutral, with a small partial being non-neutral. However, most emoji packs in misinformation are still neutral, significantly less than those in real posts. Furthermore, we analyze the correlation between the sentiment of images and text content (Appendix C.5), where 78.2% of real information contains images with the same sentiment as the text while only 34.1% of misinformation does. It further proves that misinformation and real information in MisBot are distinguishable.
Figure 14: Image distribution of misinformation and real information, including categories and sentiments. Misinformation presents a different distribution from real information.
B.3 Misinformation Detector
Figure 15: Overview of the misinformation detector, which employs multiple modality encoders to encode variance modalities and employs an MLP layer to identify misinformation.
We propose a simple misinformation detector as Figure 15 illustrates. We employ multi-modal encoders to encode content, repost, comment, image, and video. For content, we employ an encoder-based language model $\mathrm{LM}(\cdot)$ [^26] Here is the model link. to encode content:
$\boldsymbol{f}_{\textit{content}}=\mathrm{LM}(s).$
To encode repost, we employ the same language model $\mathrm{LM}(\cdot)$ to encode text-attributed node $v_{i}$ and obtain $\boldsymbol{h}_{v_{i}}^{(0)}$. We employ $L$ graph neural network layers to make each node interact:
$\boldsymbol{h}_{v_{i}}^{(\ell)}=\mathop{\mathrm{Aggr}}\limits_{\forall v_{j}\in\mathcal{N}(v_{i})}(\{\mathrm{Prop}(\boldsymbol{h}_{v_{i}}^{(\ell-1)};\boldsymbol{h}_{v_{j}}^{(\ell-1)})\}),$
where $\mathcal{N}(v_{i})$ denotes the set of neighbors of node $v_{i}$, $\mathrm{Aggr}(\cdot)$ and $\mathrm{Prop}(\cdot)$ are aggregation and propagation functions, where GCN (Kipf and Welling, 2017) is employed in practice. we finally employ the mean pooling operator as the $\mathrm{Readout}(\cdot)$ function to obtain the graph-level representation:
$\boldsymbol{f}_{\textit{repost}}=\mathrm{Readout}(\{\boldsymbol{h}_{v_{i}}^{(\ell)}\}_{v_{i}\in\mathcal{V}}).$
To encode comment, we employ the same encoding method as repost to obtain the representation of each comment graph $\mathcal{G}_{\textit{comment}}^{i}$ (Yang et al., 2023a). We then consider the average representations as the final representation:
$\boldsymbol{f}_{\textit{comment}}=\frac{1}{m}\boldsymbol{f}_{\textit{comment}}^{i},$
where $m$ is the number of comment graphs. To encode image, we employ a pre-trained swin transformer [^27] Here is the model link. (Liu et al., 2022) $\mathrm{SwinTr}(\cdot)$ to obtain the representations of each image and adopt mean pooling to obtain the final representation:
$\boldsymbol{f}_{\textit{image}}=\mathrm{mean}(\mathrm{SwinTr}(I_{i})),$
where $\mathrm{mean}(\cdot)$ denotes the meaning operator. To encode video, we sample 256 frames from each video and resize each frame into $224\times 224$. We employ pre-trained VideoMAE [^28] Here is the model link. (Tong et al., 2022) $\mathrm{VideoMAE}(\cdot)$ For each time step, we take 16 frames and set the interval to 12 frames. We could obtain:
$\boldsymbol{f}_{\textit{video}}=\mathrm{mean}(\mathrm{VideoMAE}(V_{i})).$
Finally, we concatenate them to obtain the representation of each user post:
$\boldsymbol{f}=[\boldsymbol{f}_{\textit{content}}\|\boldsymbol{f}_{\textit{repost}}\|\boldsymbol{f}_{\textit{content}}\|\boldsymbol{f}_{\textit{image}}\|\boldsymbol{f}_{\textit{video}}].$
Given an information instance $A$ and corresponding label $y$, we calculate the probability of $y$ being the correct prediction as $p(y\mid A)\propto\exp(\mathrm{MLP}(\boldsymbol{f}))$, where $\mathrm{MLP}(\cdot)$ denote an MLP classifier. We optimize this model using the cross-entropy loss and predict the most plausible label as $\arg\max_{y}p(y\mid A)$. The hyperparameter settings of the baseline are presented in Table 5 to facilitate reproduction. We conduct ten-fold cross-validation to obtain a more robust conclusion. When split folds, we do not split misinformation from the same topic (Appendix C.1) into two folds to avoid data leakage.
Hyperparameter Value Hyperparameter Value
BERT embedding dim 768 optimizer Adam
GNN layers 2 learning rate 10-4
GNN embedding dim 256 weight decay 10-5
Video embedding dim 768 dropout 0.5
Image embedding dim 768 hidden dim 256
Table 5: Hyperparameter settings of the misinformation detector.
B.4 Detector Ablation Study
We further design various variants of the misinformation detectors, removing certain components to explore which ones are essential for detection. We first remove each component except content. Then we design (i) w/o Interaction removing comment and repost; (ii) w/o Vison removing image and video; and (iii) w/o extra only containing content. For each variant, we set the remove features as $\mathbf{0}$. For example, if we remove the comment, then we set $\boldsymbol{f}_{\textit{comment}}$ as $\mathbf{0}$. We present the ablation study performance in Table 6. It illustrates that:
  • The detector without Extra modalities suffers a significant performance decline, with an accuracy drop of $17.5\%$. It is more radical, often identifying information as misinformation and achieving high recall. It proves that extra modalities provide valuable signals to identify misinformation.
  • The detector without Interaction drops to $77.3\%$ on f1-score, illustrating the effectiveness of user reaction including comments and reposts. We speculate that user interactions could provide extra evidence and signals (Grover et al., 2022) to verify the information. Meanwhile, reposts provide more evidence than comments. We assume it is related to the algorithm of social platforms, where reposted messages could be spread more widely. Thus users tend to publish verified information when reposting.
  • The detector w/o Vision only drops $2.2\%$ on f1-score, where image and video information could not provide valuable evidence. Meanwhile, video information contributes the least, with the p-value of the t-test on accuracy being $0.015$, which is not considered statistically significant. The text modalities dominate misinformation detection. We speculate that (i) annotators also consider text information when judging misinformation, introducing biases; and (ii) the pre-trained vision encoders struggle to capture signals related to identifying misinformation.
Models Accuracy F1-score Precision Recall
Vanilla $95.2_{\pm 0.6}$ $92.3_{\pm 0.8}$ $93.7_{\pm 1.7}$ $91.0_{\pm 1.0}$
w/o Comment $\mathop{93.0^{\star}_{\pm 1.4}}\limits_{2.3\%\downarrow}$ $\mathop{89.5^{\star}_{\pm 1.6}}\limits_{3.0\%\downarrow}$ $\mathop{86.1^{\star}_{\pm 3.9}}\limits_{8.1\%\downarrow}$ $\mathop{93.3^{\dagger}_{\pm 1.6}}\limits_{2.6\%\uparrow}$
w/o Repost $\mathop{89.4^{\star}_{\pm 2.0}}\limits_{6.0\%\downarrow}$ $\mathop{85.3^{\star}_{\pm 2.2}}\limits_{7.6\%\downarrow}$ $\mathop{76.8^{\star}_{\pm 4.0}}\limits_{18.0\%\downarrow}$ $\mathop{96.1^{\star}_{\pm 1.4}}\limits_{5.6\%\uparrow}$
w/o Image $\mathop{94.3^{\star}_{\pm 0.5}}\limits_{1.0\%\downarrow}$ $\mathop{90.5^{\star}_{\pm 1.0}}\limits_{1.9\%\downarrow}$ $\mathop{95.1^{\dagger}_{\pm 1.2}}\limits_{1.4\%\uparrow}$ $\mathop{86.5^{\star}_{\pm 2.0}}\limits_{4.9\%\downarrow}$
w/o Video $\mathop{95.0^{\dagger}_{\pm 0.7}}\limits_{0.2\%\downarrow}$ $\mathop{92.1^{\dagger}_{\pm 0.8}}\limits_{0.3\%\downarrow}$ $\mathop{93.0^{\dagger}_{\pm 1.9}}\limits_{0.8\%\downarrow}$ $\mathop{91.1^{\dagger}_{\pm 1.0}}\limits_{0.2\%\uparrow}$
w/o Interaction $\mathop{81.6^{\star}_{\pm 4.5}}\limits_{14.2\%\downarrow}$ $\mathop{77.3^{\star}_{\pm 4.1}}\limits_{16.2\%\downarrow}$ $\mathop{64.4^{\star}_{\pm 5.9}}\limits_{31.3\%\downarrow}$ $\mathop{97.3^{\star}_{\pm 1.2}}\limits_{7.0\%\uparrow}$
w/o Vision $\mathop{94.1^{\star}_{\pm 0.5}}\limits_{1.1\%\downarrow}$ $\mathop{90.3^{\star}_{\pm 1.0}}\limits_{2.2\%\downarrow}$ $\mathop{94.2^{\dagger}_{\pm 1.2}}\limits_{0.4\%\uparrow}$ $\mathop{86.8^{\star}_{\pm 2.1}}\limits_{4.6\%\downarrow}$
w/o Extra $\mathop{78.5^{\star}_{\pm 5.2}}\limits_{17.5\%\downarrow}$ $\mathop{74.5^{\star}_{\pm 4.3}}\limits_{19.3\%\downarrow}$ $\mathop{60.6^{\star}_{\pm 6.0}}\limits_{35.4\%\downarrow}$ $\mathop{97.3^{\star}_{\pm 1.0}}\limits_{6.9\%\uparrow}$
Table 6: Performance of the misinformation detector and variants. We report the mean and standard deviation of ten-fold cross-validation. We also report the performance changes and conduct the paired t-test with vanilla, where $\star$ denotes the p-value is less than $0.0005$ and $\dagger$ denotes otherwise. This simple misinformation detector achieves ideal performance. Misinformation and real information are distinguishable with the help of user interactions.
B.5 Expert Performance
We employ 11 social bot detectors as experts. Table 7 presents the performance and temperature of these experts. The performance of the automatic annotator is acceptable, proving the credibility of the annotations. Meanwhile, filtering in experts with an accuracy greater than 80% could improve the annotation precision. To obtain a higher precision, we set the likelihood threshold as 0.75, making sure that the annotator does not identify a genuine account as a social bot (with a precision of 97.6%).
Experts Accuracy F1-score Precision Recall Temperature
Feature-based Detectors
MLP $73.5$ $49.6$ $90.9$ $34.1$ 1.125
Random Forest $71.7$ $58.0$ $67.0$ $51.1$-
Adaboost $69.5$ $59.8$ $60.2$ $59.5$-
Content-based Detectors (BERT)
Name $74.5$ $54.3$ $86.4$ $39.6$ 1.468
Description $75.2$ $56.2$ $86.6$ $41.6$ 1.246
Posts $80.4$ $72.4$ $78.4$ $67.2$ 1.286
Content-based Detectors (DeBERTa)
Name $74.8$ $54.7$ $87.1$ $39.8$ 1.408
Description $75.2$ $58.7$ $80.9$ $46.1$ 0.972
Posts $80.6$ $73.6$ $76.9$ $70.5$ 1.129
Ensemble Detectors
BERT $83.1$ $77.3$ $79.4$ $75.4$ 1.329
DeBERTa $82.7$ $76.5$ $79.4$ $73.8$ 1.146
Annotator $85.0$ $79.5$ $83.3$ $76.1$-
All Expert $82.5$ $72.7$ $89.5$ $61.3$-
Annotator (0.75) $81.5$ $68.6$ $97.6$ $52.8$-
Table 7: The performance and temperature of the social bot detectors. The $\star$ indicates that we employ this expert in the final automatic annotator, and -- indicates that temperature scaling is not suitable for this expert. The “All Expert” denotes the ensemble of all experts. The “Annotator (0.75)” denotes that we consider an account a social bot if the likelihood is greater than 0.75.
Appendix C Details of Further Analysis
C.1 Cluster Algorithm
We cluster misinformation into different groups, where each group represents a topic or an event, based on the judgment. The main idea is that the judgments about the same event are very similar. Meanwhile, judgments about distinct events are very different. Formally, we assume the misinformation judgment set is $\{T_{i}\}_{i=1}^{N}$, where $N$ is the number of misinformation judgments. Given a specific judgment $T_{i}$, we calculate the cosine similarities of BERT [^29] Here is the model link. representations:
$s_{i,j}=\mathrm{cosine}(\mathrm{BERT}(T_{i}),\mathrm{BERT}(T_{j})).$
We sort the scores $\{s_{i,j}\}_{j=1}^{N}$ in descending order to obtain $\{s_{i,\tilde{j}}\}_{j=1}^{N}$. We then find the index $\hat{j}$ that maximize the gradient:
$\hat{j}=\arg\max_{\tilde{j}}s_{i,\tilde{j}}-s_{i,\tilde{j}+1}.$
It means judgments with a similarity score greater than $s_{i,\hat{j}}$ are very similar to $T_{i}$ and others are very distinct. Here we construct a relation from $T_{i}$ to the judgments with a similarity score greater than $s_{i,\hat{j}}$. After that, we could obtain a directed graph. We consider each strongly connected graph as a misinformation graph.
C.2 Top-ten Topics
Table 8 presents the keywords and descriptions of the top 10 topics with the highest number of misinformation items. We employ BERT [^30] Here is the model link. to obtain the representations of misinformation.
Keyword Description
Fire Disaster A place is on fire.
Dog Lost Notice Someone offers a reward of 10 million yuan to find the dog.
Import A country announced a ban on the import of another country’s coal.
Typhoon Does it feel like a disaster movie? A place is experiencing a typhoon.
Air Crash The last two minutes of a place’s air crash.
University Admission A 19-year-old freshman girl in a city fell to her death and her roommate was recommended for undergraduate study.
BBQ the woman beaten in the barbecue restaurant is dead.
Suicide The woman who jumped from a place had her home disinfected and looted.
Child Trafficking A 5-year-old son in a place was abducted near a bilingual kindergarten.
Domestic Violence The man from a province is the stepfather, and I hope the relevant departments will save this poor child.
Table 8: The keywords and descriptions of 10 topics. We translate them into English and conceal the private information.
Figure 16: The pairwise score heatmap of sematic and token levels. The values on the diagonal are significantly larger than the rest. The “Top-10” means the 10 topics with the most misinformation instances, the “Top-100” means the 100 topics with the most misinformation instances, and the “All” means all misinformation instances.
C.3 Pairwise Scores
We conduct numerical analysis to prove that misinformation in the same cluster is similar, while misinformation in different clusters is distinct. We employ semantic-level and token-level pairwise scores. Formally, we assume there are $N$ clusters (2,270 clusters), and the $i$-th cluster is represented as $\{T_{k}^{i}\}_{i=k}^{M_{i}}$, where $M_{i}$ if the number of misinformation in this cluster. Given the $i$-th cluster and $j$-th cluster, the pairwise score $s_{ij}$ is calculated as:
$s_{ij}=\frac{1}{M_{i}M_{j}}\sum_{p=1}^{M_{i}}\sum_{q=1}^{M_{j}}\mathrm{score}(T_{p}^{i},T_{q}^{j}),$
where $\mathrm{score}(\cdot,\cdot)$ is the similarity function. For semantic, we employ the cosine similarity of BERT [^31] Here is the model link. representation. For token, we employ the jieba package [^32] https://pypi.org/project/jieba/ to tokenize Chinese sentences and calculate the ROUGE-L score. Since computing pairwise ROUGE-L is time-consuming, we randomly sample 10 pieces of misinformation in each cluster.
Figure 17: Sentiment distributions in different texts. Misinformation would publish emotional content while real information would publish more neutral content. Users would publish emotional content during the information spread.
C.4 Score heatmap
Figure 16 presents the heatmap of the pairwise score, which illustrates that the values in the diagonal are much greater. It enhances our findings that: misinformation with the same topics has similar content and misinformation with different topics has distinct content.
C.5 Sentiments and Stances
To obtain the sentiments of social texts, we employ BERT trained on the EWECT dataset [^33] https://smp2020ewect.github.io/. The sentiments include neutral, happy, angry, surprised, sad, and fearful. To obtain the stances of social texts, we employ BERT trained on the STANCE dataset (Zhao et al., 2023). The stances include support, oppose, and neutral.
C.6 Sentiment Distribution
Figure 17 illustrates the distribution of sentiments in different texts. An intuitive finding is that misinformation would publish more emotional content while real news would naturally report. However, whether in misinformation or real news, public reactions are always emotional. Comments in misinformation show more anger while real news shows more happiness, both of which are more emotional than reposts. We speculate that users are inclined to comment to express emotion.
C.7 Sentiment Variation
We introduce the variation measure to calculate the degree or extent to which public sentiment changes over the news spread. Given a specific information instance and its relation comment, we first calculate the function of the proportion of comments with neutral sentiment over time $f(x)$. We then determine the time series $[x_{0},x_{1},\dots,x_{n}]$, where we set the interval as one hour. The variation is calculated as:
$v_{\Delta}=\sum_{k=1}^{n}|f(x_{k})-f(x_{k-1})|.$
C.8 Correlation Coefficient
To numerically explore the correlations between social bots and online public opinions, we calculate the following Pearson correlation coefficient:
  • The number of social bots and the number of comments with non-neural stances: 0.6661.
  • The number of social bots and the number of comments with non-neural sentiments: 0.6750.
  • The ratio of social bots and the ratio of comments with non-neural stances: 0.2040.
  • The ratio of social bots and the ratio of comments with non-neural sentiments: 0.2499.
The relatively high correlation coefficients indicate that social bots might influence public opinion.
C.9 Semantic Similarity
We explore the publishing behavior differences between social bots and genuine accounts. Here we explore whether accounts would publish similar content by introducing the semantic similarity score. Given an account with its posts in the timeline $\{T_{i}\}_{i=1}^{N}$, the semantic similarity is calculated as:
$s=\frac{1}{N^{2}}\sum_{i=1}^{N}\sum_{j=1}^{N}\mathrm{cosine}(\mathrm{BERT}(T_{i}),\mathrm{BERT}(T_{j})).$
Figure 18: The sentiments and stances of comments published by social bots. Social bots would publish polarized content, manipulating public opinions.
C.10 Social Bot Comment
We further explore the semantic information of the published content of social bots. Figure 18 illustrates the stances and sentiments of comments published by social bots. For sentiments, social bots would publish far more emotional content than neutral ones. They tend to publish happy and angry content. Namely, the large number of emotional content posted by social bots might affect public sentiment. For stances, bots would publish content with clear stances. Thus we could conclude that social bots would manipulate public opinions by publishing similar and polarized content.